### v1

# Defects and Impurity Properties of VN precipitates in ARAFM Steels: Modelling using a Universal Machine Learning Potential and Experimental Validation

Abstract
VN precipitates used to strengthen ARAFM steels for fusion applications, have been shown to undergo dissolution under high Fe ion irradiation doses of 100 dpa at dose rates of $10^{-3}$ dpa/s at 600 ${}^{o}C$. Here, point defects and solute substitutions have been studied using atom probe tomography (APT), universal machine learning interatomic potentials (uMLIPs), and density functional theory. Through a combination of transmission electron microscopy (TEM), APT, and atomic scale calculations, N-vacancies and substitutional Cr are found to be present in VN precipitates prior to irradiation. Ternary convex hulls were calculated for ten VNX (X=Cr, Fe, C, Si, Mn, W, Ta, B, S, P) systems with an uMLIP. These calculations predict Fe, P, Mn, and Si to be unstable solutes in the VN precipitate. Therefore, Fe, P, Mn and Si implantation via collision cascades is found to be a possible mechanism driving dissolution of VN.
keywords:
VN , RAFM , UMLIPs , APT
[…]
[ICL]organization=Department of Materials and Centre for Nuclear Engineering, Imperial College London, addressline=Exhibition Road, city=London, postcode=SW7 2AZ, country=UK
[…]
[UKAEA]organization=Materials Division, United Kingdom Atomic Energy Authority, addressline=Culham Campus, city=Abingdon, state=Oxfordshire, postcode=OX14 3DB, country=United Kingdom
[…]
[OX]organization=Department of Materials, University of Oxford, addressline=Parks Road, city=Oxford, state=Oxfordshire, postcode=OX1 3PH, country=United Kingdom
1 Introduction
One challenge for the commercialisation of fusion energy reactors is the development of reduced activation steels with high resistance to creep and irradiation damage that are easy to manufacture [1]. To this end, advanced reduced activation ferritic-martensitic (ARAFM) steel alloys with rocksalt-structured MX phase precipitates, such as vanadium nitride, have been proposed [2, 3, 4, 5]. However, Haley et al. recently observed that vanadium nitride precipitates underwent complete dissolution during a 100 dpa Fe+ ion irradiation [6]. ARAFM alloys are complex systems with multiple alloying elements in addition to precipitate phases. Thus, understanding the thermodynamic factors and role of irradiation damage in precipitate stability is both challenging and necessary.
Although alloy designers have often assumed MX precipitates to have idealized rocksalt stoichiometry and crystal structures, many rocksalt nitrides exhibit a strong preference for vacancy-driven non-stoichiometry [7, 8, 9, 10, 11, 12]. This preferred vacancy concentration may change if the MX phase is alloyed with other elements, either during precipitation or due to ballistic mixing. Additionally, during irradiation, the preferred vacancy concentration will be disturbed and other point defects will accumulate, which can further change the stability of the precipitate. A better understanding of the role of solutes and point defects on the stability and resilience of the VN precipitate under irradiation is needed.
We report a combined computational and experimental approach that sought to shed light on the structure of VN precipitates in ARAFM steel prior to irradiation and the types of irradiation damage that may cause dissolution. Atomic scale modelling, combined with TEM and APT results, pointed to a significant concentration of vacancies and solutes in the unirradiated precipitates. Ten V-N-X (X=Cr, Fe, C, Si, Mn, W, Ta, B, S, P) ternary systems of interest were identified as possible dissolution pathways primarily instigated by irradiation damage. Since, ternary systems represent a large undertaking for DFT calculations in terms of computational cost, we used a universal machine learning interatomic potential (UMLIP) to study these systems.
2 Methods
Samples of unirradiated ARAFM steel of the same composition and microstructure as studied under irradiation by Haley et al. were studied with TEM and APT. The lattice parameter of VN precipitates could be measured from the series of dark-field images collected by Haley et al. For each dark-field image, a corresponding diffraction pattern was captured, and the position of the objective aperture was recorded. These were compared against a simulated pattern of VN from which the lattice parameter was deduced. APT on the unirradiated ARAFM steel was collected via both voltage (20% pulse fraction, 200 kHz pulse frequency) and laser pulsing methods (40 pJ pulse energy, 200 kHz pulse frequency) at a temperature of 50 K. This data was analysed with APSuite 6.1. The precipitate compositions were measured using proxigrams created from 2.2% concentration isoconcentration surfaces with a bin size of 0.1 nm over three precipitates.
Density functional theory (DFT) calculations of point defects in VN were carried out using VASP 6 [13, 14, 15] using the Perdew, Burke, Ernzehof (PBE) exchange correlation functional [16] with k-point spacing and a planewave cut-off energy set to 0.2Å$\AA^{-1}$ and 520 eV, respectively. The projector augmented wave method (PAW) [17, 18] pseudopotentials with semicore p and s states were used for V, Fe, and Cr with Gaussian smearing of 0.1 eV for relaxation followed with a static run using the tetrahedron method with Blöchl corrections [19]. Atomic positions and unit cell parameters of all input structures were allowed to relax to a force convergence of 0.02 $/\AA$ per atom.
The thermodynamic stability of the VN and VNX systems was studied by analysing the formation energy crystal structures with point defects and clusters of point defects. Thirty intrinsic point defects were calculated with DFT: V(N) vacancies, interstitials, and anti-sites. An additional 18 DFT calculations were performed for Cr and Fe substitutions on the V sublattice and C substitutions on the N sublattice. The ideal rocksalt unit cell has a single unique site for each defect type. To vary the defect concentration, a single defect was inserted in rocksalt supercells ranging in size from 8 atoms to 256 atoms. Configurations with multiple defects of different types were not considered.
Additional calculations were performed with a universal machine learning interatomic potential (UMLIP) called Orb-v1 [20] using the atom simulation environment (ASE)[21] implemented in python. For these calculations, each structure was relaxed using the Frechet cell filter, which allowed the box and atoms to relax. The fast inertial relaxation engine (FIRE) [22] was then applied with the convergence criteria of a maximum force component of 0.05 $/\AA$ on each atom or 1000 steps.
Structures were calculated with the UMLIP as follows. First, an ideal conventional unit cell was initialized with ASE. Supercells were then enumerated using the integrated cluster expansion toolkit (ICET) [23]. The structure enumeration was broken into different series for each defect and substitution type. All possible unique structures containing the defect type were enumerated using ICET to a maximum of 14 atoms for the rocksalt structure. The Orb-v1 potential was used to calculate formation enthalpies for VN and 10 VNX systems, where X denotes the ten solute elements (Cr, Fe, C, Si, Mn, W, Ta, B, S, P). The defect types consisted of both V-N anti-sites, all three species interstitials, both V-N vacancies, and both V-N substitutions with the third species. While this was not an exhaustive search of all phases that could be included in each ternary system, it does include the rocksalt system and each element dissolved in BCC Fe, i.e. the two phases observed in experimental analysis.
Equation 1 shows an example calculation of the enthalpy of formation, $\Delta H_{f}$, for a VN structure with $n$ unit cells and interstitial solute $S$.
$\Delta H_{f}=H_{(VN)_{n}S}-(H_{(VN)_{n}}+(H_{Fe_{x}S}-H_{Fe_{X}}))$ (1)
While formation energies are typically calculated relative to the elements standard crystal structures, we have chosen to use the energy of elements as solutes in body-centered cubic (BCC) iron. This was chosen because we are most interested in the stability of these compounds when they are embedded in the steel matrix. Prior APT results had shown no V-containing precipitates other than VN before or after irradiation, which motivates the use of the energy of V and N dissolved in an Fe matrix as the reference energies. Their formation and dissolution will depend on the surrounding nearby solutes rather than pure elemental materials. The reference structures for the interstitial-type elements (C and N) consisted of 128 Fe atoms and a single interstitial atom. Both the tetrahedral and octahedral site calculations were performed, and the lowest-energy sites were selected.
3 Results
[…]
Figure 1: The lattice parameters calculated using VASP for defects in rocksalt VN as a function of concentration are compared to the experimental value (unknown stoichiometry) from TEM highlighted with a horizontal line. The effect of vacancies, interstitial, and antisites are compared in a). b) shows a convex hull calculated with over 800 enumerated structures calculated with the Orb-v1 UMLIP. The red line indicates the convex hull with diatomic N and BCC V as reference states, while the blue line is the convex hull where V and N as dilute solutes in Fe are used as reference states.
The Orb machine learning potential was tested against DFT to ensure accuracy (Table 1). The Orb potential aligns well with the DFT values for stable structures and vacancies. However, the machine learning potential does not capture the high-energy unstable structures or structures containing interstitial-type defects well. This is likely due to bias from the training data, which includes vacancy-like defects but limited interstitial-type and high-energy structures.
Table 1: The root mean squared error (RMSE) values for UMLIP values against the DFT values for point defects in VN calculated over 52 structures.
[…]
Overall 0.36 0.28 0.52 0.29
V-N anti site 0.69 0.26 0.73 0
N-V anti site 0.30 0.25 0.37 0.37
V vacancies 0.14 0.07 0.09 0.03
N vacancies 0.14 0.02 0.03 0.03
V interstitials 0.40 0.39 0.48 0.42
N interstitials 0.17 0.38 0.22 0.11
The UMLIP calculated convex hull is shown in Fig 1 b). This convex hull is in good agreement with a previously reported VN convex hull[10] in that it reproduces two key features: non-stoichiometric N-vacancy-driven rocksalt structure are predicted as ground states, and these structures have highly ordered layered-vacancy structures. However, the effect of the chosen reference states plays a role in the relative stability of N vacancies when the VN is considered to be part of the wider system with the solute reference states. When the system is referenced to V and N as dilute solutes in Fe, the hull is pulled below the ground states with higher N vacancy concentration. Structures with lower N vacancy concentrations remain on the hull. These calculations combined with the observed contracted lattice parameter suggest that N vacancies may be present in the VN precipitates in significant concentrations prior to irradiation.
Figure 2: a) The concentration profiles of decomposed V, N, and Cr through the interface of a 2.2% isoconcentration surface of an APT reconstruction averaged over 3 precipitates. b) The concentration profiles of decomposed C, Si, W and Mn through the interface of a 2.2% isoconcentration surface of an APT reconstruction. For both plots, the distance is the distance from the interface of the isosurface, with positive values inside the precipitate and 0.0 nm being the interface marked by a vertical dashed line.
Figure 2 shows the proxigrams of V, N, Cr, Si, W, Mn, and C across all interfaces from RAFM steel matrix and VN precipitate in a unirradiated sample. Cr, Si, W, Mn, and C were all present inside the precipitates with Cr and Si concentrations increasing towards the centre of the precipitates. The vanadium concentration is approximately double that of the measured nitrogen concentration, Figure 2 a), though this is likely exaggerated due to an under-measurement of N by APT. Some N type vacancies may be present as previously discussed, though the concentration is likely less pronounced than this measurement suggests.
DFT calculations of the lattice parameters as a function solute concentration are compared to the observed lattice parameters in Fig. 3 a). Substitutions of carbon on the nitrogen sublattice increase the lattice parameters while both Fe and Cr substitutions on the vanadium sublattice are predicted to decrease the lattice parameter. Although substitutions change the lattice parameter, the effect is minimal with changes being less than 0.2%. Thus, Cr and Fe substitutions may slightly alter the lattice parameter but are not a significant drivers of the lattice contraction measured here.
Figure 3: The effect of substitutional C, Fe, and Cr on lattice parameter as calculated by DFT are plotted in a) and compared to an experimental value from a single measurement at an unknown stoichiometry (the dashed horizontal line). Three ternary phase diagrams (b) V-N-C, c) V-N-Cr, d) V-N-Fe) were calculated using Orb-v1 where all points shown are on the convex hull.
To better understand the effect of these solutes in VN before irradiation, and the effects of compositional changes during irradiation, the convex hulls (Fig. 3) of three V-N-X ternary systems were calculated with DFT where X= Cr, Fe and C. In the ternary convex hull plots, the V-N-Cr system (Fig. 3 c) shows stable Cr substitutions on the V sites, which may explain the presence of Cr in the VN precipitates as found using APT. The VN precipitates show limited C content from the APT data, as shown in Fig. 2. This does not align with the ML potential enumerated convex hull shown in Fig. 3 b), where C is found on N sites. However, it is known that C content is underestimated in APT due to its tendency for high multiple detector events [25].
The presence of Fe in VN is particularly important when considering irradiation damage in Fe-based RAFM steels. Under high-energy collision cascades, Fe atoms will be ballistically implanted and point defects will accumulate at the interface between the Fe and the VN. Therefore, because the FeN rocksalt is unstable, Fe substitution in VN could destabilise and contribute to precipitate dissolution. Unfortunately, the Fe concentration in VN is difficult to determine using APT due to the field evaporation-induced artefacts. In addition, Si was found at higher concentrations in VN with APT, and this may also be an artefact of APT due to the ease of migration of Si [26]. Fe had no stable structures on the hull, as shown in Fig. 3 d). This indicates Fe substitutions or Fe interstitial accumulation may act to destabilize the VN.
Seven more ternary systems have been investigated using the Orb interatomic potential: V-N with P, S, Mn, W, Ta, Si and B Fig 4. These additional elements have been investigated because they are present in ARAFM steel [6] or, in the case of B, in boron reduced activation ferritic-martensitic steel (BRAFM) steel [27]. Each ternary plot includes the V-N binary hull already discussed on the right most edge as well as elemental reference structures at each vertex. Additional points plotted on the ternary diagram indicate additional stable structures in the ternary system. Our study did not yield any additional rocksalt-derived structures on the convex hull for the Si and P systems (Fig. 4 e) and c)). However, the remaining S, W, Ta, Mn, and B all contain stable solute-containing structures on the hull, Fig. 4 a), b), d), f) and g). As expected, the smaller radii elements of B (Fig. 4 a) and S Fig. 4 d) occupy the N sites, while the larger W and Ta occupy the V sites. Interstitial solutes were predicted to be unstable in all of the ternary systems studied.
Figure 4: Seven ternary systems are presented as ternary phase diagrams, calculated using Orb-v1, where all points are on the convex hull. a) V-N-B, b) V-N-Mn, c) V-N-P, d) V-N-S, e) V-N-Si, f) V-N-Ta, g) V-N-W.
4 Discussion
APT results show that VN precipitates may not be modelled appropriately as only a V-N system and that ternary or more complex systems may be required. However, DFT, whilst commonly used for its accuracy, is too computationally expensive to study such systems. Here, the Orb UMLIP was employed to explore a larger series of systems that cannot easily be explored with DFT alone, and it can do so capably. The UMLIP was able to approach DFT-level results for stable VN structures [10], including finding previously reported stable vacancy layers.
[…]
Both this work and prior work in the literature[28] indicate that VN accommodate N vacancies, with highly ordered layered vacancy structures as ground states. However, irradiation damage at 10s of dpa is expected to destroy long-range ordering in these precipitates. The particular temperature and fluence experienced by the material could result in a dynamic equilibrium by which irradiation destroys ordered structures generating randomly ordered vacancies, and a thermodynamic driving force drives diffusion to reorder the structure into vacancy layers.
In addition, the ternary systems studied suggest possible mechanisms of dissolution of VN. Due to irradiation damage, Fe is likely to be present in the VN precipitates, either via ballistic implantation, or by vacancy-driven migration. However, our calculations (Fig. 3) indicate that Fe may act to destabilise the VN.
5 Conclusions
A combined experimental and computational approach have led to both a better understanding of the VN precipitates in ARAFM steel and the effect of common irradiation defect on precipitate stability. These findings can be summarized as follows.
  • 1.
    A contracted lattice parameter measurement and APT results show that Cr, along, with other solutes, are present in these precipitates, suggesting the VN precipitates should not be modelled as an ideal binary rocksalt phase.
  • 2.
    The small experimentally measured lattice parameter for VN precipitates can be at least partially explained by a combination of N-vacancies and solutes. Stoichiometric vacancies on both sublattices may also contribute to the change in lattice parameter, though this was not explored in this study.
  • 3.
    UMLIP and DFT results both predict sub-stoichiometric ground states in the VN system with ordered vacancy layers favoured over disordered point vacancies.
  • 4.
    UMLIP calculations predict Fe substitutions to be unstable in VN.
  • 5.
    Switching from bulk to local reference states changed the apparent stability of vacancies in VN. This highlights how a calculated binary phase diagram may lead to inaccurate expectations when predicting phase stability of precipitates in more complex alloy systems, such as MX precipitates in ARAFM steel. Reconsidering the reference states is a simple way to gauge phase stability in these alloys. Further calculations of the phase stabliity of other ARAFM steel MX phase precipitates with respect to local reference states are needed.
Acknowledgements
We would like to acknowledge computational resources and support provided by the Imperial College Research Computing Service (http://doi.org/10.14469/hpc/2232). We are grateful to the UK Materials and Molecular Modelling Hub for computational resources, which is partially funded by EPSRC (EP/T022213/1, EP/W032260/1 and EP/P020194/1). DNM and TM would like to acknowledge the EUROfusion support for providing access to Marconi-Fusion HPC facility in generation of the DFT data used for this paper. This work has been funded by the NEUtron iRradiatiOn of advaNced stEels (NEURONE) programme via Fusion Futures. As announced by the UK Government in October 2023, Fusion Futures aims to provide holistic support for the development of the fusion sector.