### v2

# Defects and Impurity Properties of VN Precipitates in ARAFM Steels: Modelling using a Universal Machine Learning Potential and Experimental Validation

Abstract
[…]
keywords:
Vanadium nitride , Reduced activation ferritic martensitic steel , Universal machine learning potential , Atom probe tomography
journal: Journal of Nuclear Materials
[…]
[ICL]organization=Department of Materials and Centre for Nuclear Engineering, Imperial College London, addressline=Exhibition Road, city=London, postcode=SW7 2AZ, country=UK
[…]
[UKAEA]organization=Materials Division, United Kingdom Atomic Energy Authority, addressline=Culham Campus, city=Abingdon, state=Oxfordshire, postcode=OX14 3DB, country=United Kingdom
[…]
[OX]organization=Department of Materials, University of Oxford, addressline=Parks Road, city=Oxford, state=Oxfordshire, postcode=OX1 3PH, country=United Kingdom
{highlights}
APT results show that Cr, along, with other solutes, are present in these precipitates.
The small experimentally measured lattice parameter for VN precipitates can be at least partially explained by a combination of N-vacancies and solutes.
NEP and DFT results both predict sub-stoichiometric ground states in the VN system with ordered vacancy layers favoured.
Switching from bulk to local reference states changed the apparent stability of vacancies in VN.
Monte Carlo and cascades confirm the presence of ordered vacancies at high temperature and their role in mitigating irradiation damage
1 Introduction
One challenge for the commercialisation of fusion energy reactors is the development of reduced activation steels with high resistance to creep and irradiation damage that are easy to manufacture [30]. To this end, advanced reduced activation ferritic-martensitic (ARAFM) steel alloys with rocksalt-structured MX phase precipitates, such as vanadium nitride (VN), have been proposed [10, 27, 17, 25]. However, Haley et al. recently observed that VN precipitates underwent complete dissolution during a 100 dpa Fe+ ion irradiation [16]. ARAFM alloys are complex systems with multiple alloying elements in addition to precipitate phases. Thus, understanding the thermodynamic factors and role of irradiation damage in precipitate stability is both challenging and necessary.
Although alloy designers have often assumed MX precipitates to have idealized rocksalt stoichiometry and crystal structures, many rocksalt nitrides exhibit a strong preference for vacancy-driven non-stoichiometry at 0 K[2, 3, 23, 32, 14, 13]. This preferred vacancy concentration may change if the MX phase is alloyed with other elements, either during precipitation or due to ballistic mixing. Additionally, during irradiation, the preferred vacancy concentration will be disturbed and other point defects will accumulate, which can further change the stability of the precipitate. A better understanding of the role of solutes and point defects on the stability and resilience of the VN precipitate under irradiation and different temperatures is needed.
We report a combined computational and experimental approach that sought to shed light on the structure of VN precipitates in ARAFM steel prior to irradiation and the types of irradiation damage that may cause dissolution. Atomic scale modelling, combined with transmission electron microscopy (TEM) and APT results, pointed to a significant concentration of vacancies and solutes in the unirradiated precipitates. Due to the high temperature operating conditions of ARAFM steels, we performed Monte Carlo and molecular dynamics at temperature using a finetuned the universal neuroevolution potential (NEP) [12, 26]. The finetuned NEP potential was then used to study the effects of irradiation damage and temperature on stability of the rocksalt V-N-X (X=Cr, Fe, C, Si, Mn, W, Ta, B, S, P) system.
2 Methods
Samples of unirradiated ARAFM steel of the same composition and microstructure, as studied under irradiation by Haley et al, were studied with TEM and APT. The lattice parameter of VN precipitates could be measured from the series of dark-field images collected by Haley et al. For each dark-field image, a corresponding diffraction pattern was captured, and the position of the objective aperture was recorded. These were compared against a simulated pattern of VN, from which the lattice parameter was deduced. APT on the unirradiated ARAFM steel was collected via both voltage (20% pulse fraction, 200 kHz pulse frequency) and laser pulsing methods (40 pJ pulse energy, 200 kHz pulse frequency) at a temperature of 50 K. This data was analysed with APSuite 6.1. The precipitate compositions were measured using proxigrams created from 2.2% concentration isoconcentration surfaces with a bin size of 0.1 nm over three precipitates.
[…]
The thermodynamic stability of the VN and VNX systems was studied by analysing the formation energy crystal structures with point defects and clusters of point defects. Thirty intrinsic point defects were calculated with DFT: V(N) vacancies, interstitials, and anti-sites. An additional 18 DFT calculations were performed for Cr and Fe substitutions on the V sublattice and C substitutions on the N sublattice. The ideal rocksalt unit cell has a single unique site for each defect type. To vary the defect concentration, a single defect was inserted in rocksalt supercells ranging in size from 8 to 256 atoms. Configurations with multiple defects of different types were not considered as the primary objective was to isolate the effect of each defect. This approach necessarily limits the analysis to non-interacting defects.
To ensure consistency in the training data, the Materials Project trajectory DFT parameters [18] were used for training data generation, as the importance of maintaining consistency has been demonstrated by Mazitov et al. [28]. The training data consisted of 2797 structures, as detailed in Table 1, containing VN, VN doped with solutes, body-centered cubic Fe with solutes and ordered structures from the materials project database [18]. The data was split into a training and testing set at a 80:20 ratio. In total 11 solutes were included; Fe, B, C, Cr, Mn, P, S, Si, Ta, Ti and W. The inclusion of 11 solutes was to enable the subsequent modelling of true chemical complexity found in ARAFM steel. However, such a large element count will necessarily lower benchmark accuracy against DFT data. These structures were generated using NepTrainKit [9] and the Integrated cluster expansion toolkit (ICET). The non-ICET enumerated structures had an atomic position rattle of 0.05 Å and a random lattice scaling of 1%. All non-enumeration structures were subsampled down by a factor of five using farthest point sampling to ensure training efficiency [11]. The enumeration structures were generated using ICET [1] to enumerate all possible configurations of defects and substitutions in the V-N-X system up to 14 atoms for VN and 11 for V-N-X with X denoting one of the 11 solute elements.
Table 1: Summary of DFT training structures
System Number of structures
VN with solutes
   Structure enumeration (up to 12 atoms) 1569
   Randomly doped VN (32 atoms) 175
VN
   Structure enumeration (up to 14 atoms) 867
   Large VN structures (64 to 96 atoms) 54
   Large VN structures with vacancies (31 to 95 atoms) 22
Other
   Solute reference states in Fe (128 to 129 atoms) 13
   Ordered phases (up to 126 atoms) 97
Total 2797
The NEP potential was fine-tuned from the NEP89 foundational model [26]. The potential was fine-tuned with the following training parameters
$\lambda_{1}=\lambda_{2}=0.2,$ (1)
$\lambda_{E}=10,$ (2)
$\lambda_{F}=\lambda_{V}=1,$ (3)
where $\lambda_{1}$ and $\lambda_{2}$ are the $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ norm regularisations respectively. $\lambda_{E}$, $\lambda_{F}$ and $\lambda_{V}$ are the energy, forces and virials training weights respectively. In addition, the training was run for 5000 generations with a population size of 20 and a batch size of 2244, chosen to match the entire training dataset size. The model performance is presented in Table 2. The root mean squared error (RMSE) values improved during finetuning and reached comparable errors to large message passing neural networks, such as orb-v3 [33], but with a substantially lower inference computational cost of approximately three orders of magnitude. The testing dataset parity plots for energy, forces, stresses and virials are included in the supplementary information in fig. 9. This reduced computational cost was prioritised over only minimal errors to be capable of performing molecular dynamics of large systems. Crucially the NEP potential after finetuning was able capture the same vacancy ordered structures on the convex hull as DFT indicating its suitability.
Table 2: Model performance metrics
Metric Error
RMSE test energy 0.068 eV
RMSE test forces 0.577 eV/Å
RMSE test virials 0.263 eV/atom
RMSE VN hull distance 0.242 eV/atom
RMSE VN hull distance 0.093 eV/atom
   (within 0.2 eV)
[…]
Structures were calculated with the NEP as follows. First, an ideal conventional unit cell was initialized with ASE. Supercells were then enumerated using ICET [1]. The structure enumeration was broken into different series for each defect and substitution type. All possible unique structures containing the defect type were enumerated using ICET to a maximum of 14 atoms for the rocksalt structure. The NEP potential was used to calculate formation enthalpies for defect in VN. The defect types consisted of both V-N anti-sites, both V and N interstitials and both V-N vacancies. While this was not an exhaustive search of all phases that could be included in the system, it does include the unique single type defect in the VN rocksalt system and each element dissolved in BCC Fe, i.e. the two phases observed in experimental analysis.
Equation 4 shows an example calculation of the enthalpy of formation, $\Delta H_{f}$, for a VN structure with $n$ unit cells and interstitial solute $S$.
$\Delta H_{f}=H_{(VN)_{n}S}-(H_{(VN)_{n}}+(H_{Fe_{x}S}-H_{Fe_{X}}))$ (4)
While formation energies are typically calculated relative to the elements standard crystal structures, we have chosen to use the energy of elements as solutes in (BCC) iron. This was chosen because we are most interested in the stability of these compounds when they are embedded in the steel matrix. Prior APT results had shown no V-containing precipitates other than VN before or after irradiation, which motivates the use of the energy of V and N dissolved in an Fe matrix as the reference energies. Their formation and dissolution will depend on the surrounding nearby solutes rather than pure elemental materials. The reference structures for the interstitial-type elements (C and N) consisted of 128 Fe atoms and a single interstitial atom. Both the tetrahedral and octahedral site calculations were performed, and the lowest-energy sites were selected.
To investigate the ordering of vacancies at the high temperatures targeted by ARAFM steel of around 900 K an on-lattice Monte Carlo approach was taken to address the 0 K limitation of DFT. A cubic supercells containing 2744 atoms were used with different initial compositions of VN with solutes and vacancies. The vacancies are represented as a unique species labelled X. A Monte Carlo algorithm was then used to find a representative low energy configuration for each of the given conditions. The simulations were carried out on-lattice to enable the swapping of vacancies, which remains difficult for off-lattice simulations.
The Monte Carlo algorithm used is as follows. Where $E_{i}$ is the energy before a swap, $E_{f}$ is the energy after a swap, $k_{B}$ is the Boltzmann constant and $T$ is temperature.
$\Delta E=E_{f}-E_{i}$ (5)
[…]
To carry out the Monte Carlo simulations the software package Clease [8] was used with the finetuned NEP potential. The simulations were run for a total of 70,000 attempted swaps for 3 conditions; VN at 933 K with different initial vacancy concentrations, VN with 25% vacancies on the N sublattice at temperatures between 0 and 2000 K and VN with solutes and vacancies representative of the APT concentration of 5% vacancy and 5% C substitutions on the N sublattice and 12% Cr, 1% Mn, 1% W, 2% Si and 5% Fe on the V sublattice.
To quantify the vacancy ordering, the short range order (SRO) parameter was quantified for each neighbour shell, k, from 1 to 10. The SRO was calculated by
$\alpha_{ij}^{(k)}=1-\frac{n_{ij}^{(k)}}{2c_{j}}$ (7)
where $\alpha_{ij}^{(k)}$ is the SRO parameter between elements $i$ and $j$ for neighbour shell $k$, $n_{ij}^{(k)}$ is the number of atoms of type $j$ neighbouring atoms of type $i$ in shell $k$, $c_{j}$ is the concentration of atoms of type $j$ in the bulk, and the factor of 2 scales the SRO to one sublattice.
The final structures from the following Monte Carlo runs at 933 K; (5%, 14%, 26%, 50%) vacancy concentration and the APT representative concentration were used as the starting configurations for primary knock-on atom (PKA) cascades. The PKA cascades were carried out three times for each of the [100], [110] and [111] directions and for each starting structure. The PKA energy was set to 1200 eV and the temperature of the simulation was set to 900 K. The structure were first relaxed using the FIRE algorithm for 2000 steps and then thermalised using the stochastic rescaling NPT ensemble [4] for between 5 ps to 14 ps. The thermalisation time was varied between each run of the same structure to result in a different random seed for the cascade simulation. After thermalisation, the central N atom was selected and given a kinetic energy of 1200 eV. The NVE ensemble was used in the central region and a Nose-Hoover chain (NHC) thermostat was set to a 10 Å thick boundary at the outer edge of the supercell. The NHC thermostat was given a target temperature of 900 K with a temperature coupling constant of 200 K. The timestep was set at 1 fs with a maximum displacement per step of 0.015 Å.
To quantify the damage following the cascades the change in stored energy was calculated. This was used instead of the defect count that would be unreliable in this case due to the presence of vacancy defects in the initial structures. The stored energy, $\Delta E_{\mathrm{stored}}$, is given by the relaxed energy after the cascade, $E_{\mathrm{final}}$ minus the relaxed energy before the cascade $E_{\mathrm{initial}}$, or simply
$\Delta E_{\mathrm{stored}}=E_{\mathrm{final}}-E_{\mathrm{initial}}.$ (8)
3 Results
The experimental lattice parameter for the VN precipitates as measured with TEM was $4.035\r${A}, a smaller measurement than the previously reported ($4.1287\r${A}) [24] and much smaller than the calculated lattice parameter for the pristine VN rocksalt structure. This measurement is compared to the lattice parameter of VN as a function of the concentration of intrinsic point defects (vacancies, antisites, and interstitials) as calculated with DFT in Figure 1a). Vacancies on both the vanadium or nitrogen sublattices both lead to a reduction in the lattice parameter (Figure 1a) of up to 2%, whereas the anti-sites and interstitials increase the lattice parameter by 4% and 8% respectively.
Figure 1: The lattice parameters calculated using VASP for defects in rocksalt VN as a function of concentration are compared to the experimental value (unknown stoichiometry) from TEM highlighted with a horizontal line. The effect of vacancies, interstitial, and antisites are compared in a). The effect of substitutional C, Fe, and Cr on lattice parameter as calculated by DFT are plotted in b) and compared to an experimental value from a single measurement at an unknown stoichiometry (the dashed horizontal line).
The NEP calculated convex hull is shown in Fig 2. This convex hull is in good agreement with a previously reported VN convex hull[32] in that it reproduces two key features: non-stoichiometric N-vacancy-driven rocksalt structures are predicted as ground states, and these structures have highly ordered layered-vacancy structures. However, the effect of the chosen reference states plays a role in the relative stability of N vacancies when the VN is considered to be part of the wider system with the solute reference states. When the system is referenced to V and N as dilute solutes in Fe, the hull is pulled below the ground states with higher N vacancy concentration. Structures with lower N vacancy concentrations remain on the hull. These calculations combined with the observed contracted lattice parameter suggest that N vacancies may be present in the VN precipitates in significant concentrations prior to irradiation.
Figure 2: The convex hull calculated with 6519 enumerated structures calculated with the finetuned NEP potential. The red line indicates the convex hull with molecular N and BCC V as reference states, while the blue line is the convex hull where V and N as dilute solutes in Fe are used as reference states.
Figure 3 shows the proxigrams of V, N, Cr, Si, W, Mn, and C across all interfaces from RAFM steel matrix and VN precipitate in a unirradiated sample. Cr, Si, W, Mn, and C were all present inside the precipitates with Cr and Si concentrations increasing towards the centre of the precipitates. The vanadium concentration is approximately double that of the measured nitrogen concentration, Figure 3 a), though this is likely exaggerated due to an under-measurement of N by APT. Some N type vacancies may be present as previously discussed, though the concentration is likely less pronounced than this measurement suggests.
DFT calculations of the lattice parameters as a function solute concentration are compared to the observed lattice parameters in Fig. 1 b). Substitutions of carbon on the nitrogen sublattice increase the lattice parameters while both Fe and Cr substitutions on the vanadium sublattice are predicted to decrease the lattice parameter. Although substitutions change the lattice parameter, the effect is minimal with changes being less than 0.2%. Thus, Cr and Fe substitutions may slightly alter the lattice parameter but neither are a significant driver of the lattice contraction measured here.
Figure 3: a) The concentration profiles of decomposed V, N, and Cr through the interface of a 2.2% isoconcentration surface of an APT reconstruction averaged over 3 precipitates. b) The concentration profiles of decomposed C, Si, W and Mn through the interface of a 2.2% isoconcentration surface of an APT reconstruction. For both plots, the distance is the distance from the interface of the isosurface, with positive values inside the precipitate and 0.0 nm being the interface marked by a vertical dashed line.
Ordered vacancy layers have been confirmed in this work at zero temperature and with respect to solute reference state, as shown in the VN convex hull in Fig. 2. However, to investigate the effect of temperature on-lattice Monte Carlo simulations were conducted. The SRO between vacancy-vacancy at 933 K in the first to tenth nearest neighbour shells (the sites are illustrated in Fig. 4 c)) as a function of vacancy concentration was found to trend to zero and random ordering with increasing vacancy concentrations, as shown in Fig. 4 a). In addition, a high SRO parameter can be seen for the shell numbers 2 and 4 highlighting less than random occupations. The sixth and seventh neighbour shells are found to have the lowest SRO parameters, although the ordering decreases with increasing vacancy concentration. The effect of temperature was subsequently investigated with a fixed vacancy concentration of 25%. As with the vacancy concentration, increasing temperature was found to decrease ordering, as shown in Fig. 4 b). It is noted that significant ordering still exists at 933 K for shells 2 and 6. Shell 4 exhibits unusual behaviour likely a result of fluctuations between many configurations close in energy. The APT representative composition was also investigated at 933 K and the SRO as function of shell number is shown in Fig. 4 d). The Cr, Mn, and Si show the most ordering with all solutes having weaker ordering at larger distances.
Figure 4: The ordering present taken from the final states from on-lattice Monte Carlo simulations. a) shows the SRO parameters as a function of initial vacancies (X) on the N sublattice at a temperature of 933 K. b) shows the SRO parameters as a function of temperature at a vacancy concentration of 25%. c) an illustration of nearest neighbour shells of N atoms in a slice in the [111] plane viewed down the [111] direction. d) shows the SRO parameters as a nearest neighbour shells including solutes at a temperature of 933 K.
The vacancy ordering behaviour showed stark differences with VN and VN with the added solutes. The vacancies on the N sublattice has been shown to order along the [111] plane. However, with the introduction of solute atoms, most notably large atoms such as W, this ordering is broken down and the formation of clustering like behaviour was observed, as shown in Fig. 5. In particular, in Fig. 5 a) and b) the vacancies form into vacancy-W-Si-Mn clusters. This is likely partially because the presence of vacancies mitigating the strain induced from the large W atoms.
Figure 5: The on-lattice Monte Carlo atomic configurations taken from the final state of a 5% vacancy concentration at 933 K with and without solutes. a) and c) show the vacancies in $\mathrm{VN}_{0.95}$ condition coloured by species, a), or by local vacancy concentration (within 3 Å). b) and d) show the vacancies in $\mathrm{VN}_{0.95}$ containing solutes condition coloured by species, a), or by local vacancy concentration (within 3 Å).
To understand the dynamics of irradiation damage, PKA collision cascades were performed on five conditions; VN, VN with a (5%, 14%, 26%, 50%) pre-existing vacancy concentration and the APT representative concentration. All starting structures were taken from the final state of the Monte Carlo simulations. As a result structures containing vacancies started with ordered vacancies, and the APT representative structure contained vacancy-solute clusters. The lower initial vacancy concentration conditions of VN0.86, VN0.95 and VN0.74 had a lower increase in stored energy when compared with the pristine VN, as shown in Fig. 6 a). The higher vacancy concentration VN0.5 was found to have a larger increase in stored energy over the pristine case. Finally, the APT composition had an apparent protective effect due to the presence of 5% vacancies and an apparent detrimental effect due to the addition of solutes. The two competing effects result in a comparable stored energy to the pristine condition. The Frenkel pair count over time, given by Fig. 6 b), shows a clear relationship between vacancy number and final number of Frenkel pairs. However, this is caused by the measurement including the movement of defect as opposed the their creation. Therefore, the stored energy is a more applicable metric that include the increase in energy from both the disordering of defects as well as the creation of new defects.
Figure 6: The damage introduced from collision cascades. a) showing the change in stored energy for each condition. b) The Frenkel pair count over time for each condition calculated from the initial frame starting condition.
4 Discussion
APT results show that VN precipitates in these ARAFM may not be modelled appropriately as only a V-N system and that ternary or more complex systems may be required. However, DFT, whilst commonly used for its accuracy, is too computationally expensive to study such systems. Here, the NEP89 potential was finetuned and employed to explore a larger systems that cannot easily be explored with DFT alone, and it can do so capably. The potential was able to replicate DFT-level results for stable VN structures [32], including finding previously reported stable vacancy layers.
Experimentally, the pre-irradiation vacancy and solute concentration for the VN precipitates is difficult to measure due to the small precipitate size and the limitations of APT. The point defect calculations for VN show a preference for vacancy-driven non-stoichiometry with a thermodynamic preference for layered vacancy structures proposed by [14, 13]. However, when accounting for the local environment by adjusting the reference states the preference for N vacancies was weaker than previously reported reducing the range of stable stoichiometries present, highlighting a need to reconsider MX-type precipitates in the context of the surrounding steel matrix. The low vacancy concentration structures (5% to 50%) were still present on the convex hull. Including N vacancies and Cr-V substitutions may partially explain the small experimental lattice parameter of 4.035 Å, although the Cr had little effect.
Both this work and prior work in the literature[31] indicate that VN can accommodate N vacancies, with highly ordered layered vacancy structures as ground states. However, irradiation damage at 10s of dpa is expected to destroy long-range ordering in these precipitates. The particular temperature and fluence experienced by the material could result in a dynamic equilibrium by which irradiation destroys ordered structures generating randomly ordered vacancies, and a thermodynamic driving force drives diffusion to reorder the structure into vacancy layers. It is shown here that there remains ordering of vacancies at the target operating temperature of over 800 K, albeit the ordering is reduced. In addition, the effect of single cascades suggests that low to moderate vacancy concentrations of (5% to 50%) may be more robust to irradiation damage on account of deposited incoming energy leading to the reordering of defects in addition to creating them. Lastly, the effect of solutes seems to play a major role in irradiation damage of VN, and as such caution should considered when such solutes are excluded. In particular, solutes act to disrupt the vacancy ordering on account of introducing strain into the lattice. Such strain leads to the formation of solute plus vacancy clusters and is shown with cascades (Fig. 6) to increase the stored energy relative to the 5% vacancy condition. Whilst the lattice Monte Carlo model showed vacancy-solute cluster formation, it is important to recognise the model does not account for kinetics. It is known that cluster composition and formation depends partially on the diffusion rates of solutes and defects. In particular, it is found here that W is a key component of solute clustering in VN, although given its large atomic radii it is expected to diffuse slowly. A model accounting for both the energetics and kinetics may not find the same vacancy behaviour.
5 Conclusions
A combined experimental and computational approach have led to both a better understanding of the VN precipitates in ARAFM steel and the effect of common irradiation defects on precipitate stability. These findings can be summarized as follows.
  • 1.
    A contracted lattice parameter measurement and APT results show that Cr, along, with other solutes, are present in these precipitates, suggesting the VN precipitates should not be modelled as an ideal binary rocksalt phase.
  • 2.
    The small experimentally measured lattice parameter for VN precipitates can be at least partially explained by a combination of N-vacancies and solutes. Stoichiometric vacancies on both sublattices may also contribute to the change in lattice parameter, though this was not explored in this study.
  • 3.
    NEP and DFT results both predict sub-stoichiometric ground states in the VN system with ordered vacancy layers favoured over disordered point vacancies.
  • 4.
    Switching from bulk to local reference states changed the apparent stability of vacancies in VN. This highlights how a calculated binary phase diagram may lead to inaccurate expectations when predicting phase stability of precipitates in more complex alloy systems, such as MX precipitates in ARAFM steel. Reconsidering the reference states is a simple way to gauge phase stability in these alloys. Further calculations of the phase stability of other ARAFM steel MX phase precipitates with respect to local reference states are needed.
  • 5.
    Lattice based Monte Carlo confirmed the presence of vacancy ordering at operational temperatures of 800-900+ K in ARAFM steel.
  • 6.
    PKA collision cascades showed the importance of vacancies and solutes in stored defect energy introduced from cascades. In particular, the 5, 14 and 26% vacancy concentrations mitigated damage, whilst the addition of solutes reversed this for the 5% solute case tested.
6 CRediT authorship contribution statement
R. S. Stroud: Conceptualization, data curation, investigation, formal analysis, writing—original draft, and visualization. C. Reynolds: Writing—review and editing and supervision. T. Melichar: Data curation, investigation, formal analysis and visualization. J. Haley: Investigation and writing—review and editing. M. Carter: investigation. M. Moody: Funding acquisition. C. Hardie: Supervision, writing—review and editing. D. Bowden: Supervision. D. Nguyen-Manh: Conceptualization, formal analysis, writing—review and editing and supervision. M. R. Wenman: Conceptualization, formal analysis, writing—review and editing, supervision, and funding acquisition
Acknowledgements
We would like to acknowledge computational resources and support provided by the Imperial College Research Computing Service (http://doi.org/10.14469/hpc/2232). We are grateful to the UK Materials and Molecular Modelling Hub for computational resources, which is partially funded by EPSRC (EP/T022213/1, EP/W032260/1 and EP/P020194/1). DNM and TM would like to acknowledge the EUROfusion support for providing access to Marconi-Fusion HPC facility in generation of the DFT data used for this paper. This work has been funded by the NEUtron iRradiatiOn of advaNced stEels (NEURONE) programme via Fusion Futures. As announced by the UK Government in October 2023, Fusion Futures aims to provide holistic support for the development of the fusion sector.
Supplementary Information
Figure 7: Density functional theory convergence testing for kpoint spacing at 520 eV energy cutoff.
Figure 8: Convergence testing for energy cutoff at a kpoint spacing of 0.2 Å-1.
Figure 9: The testing dataset parity plots for the finetuned NEP89 potential against the DFT values for a) energy, b) force, c) stress, and d) virial.